The numbers 5, 10, 15, and 20 are written on slips of paper, which are then placed in a hat. Two slips are randomly selected with replacement. Find the mean of the samples where the first is smaller than the second.

1
Expert's answer
2022-02-22T17:09:29-0500

Solution:

Given numbers = 5, 10, 15, and 20

Samples of size 2 where the first is smaller than the second= {(5,10),(5,15),(5,20),(10,15),(10,20),(15,20)}


Mean of samples"=\dfrac{\frac{5+10}2+\frac{5+15}2+\frac{5+20}2+\frac{15+10}2+\frac{20+10}2+\frac{15+20}2}{6}"

"=\dfrac{7.5+10+12.5+12.5+15+17.5}6
\\=\dfrac{75}6
\\=12.5"